Monday Morning Meeting April 21st

  • USA Toll Free Number: 866-901-8266
  • USA Toll Number: +1-203-566-3863


  • Socorro: Rob, Lindsey, Susan, George, Tak.
  • CV: Andy, Mark, Dave, Darrell, Juergen.
  • Garching: N/A
  • SCO:

  • (Kumar and Sandra on leave. Dirk Jim and Jeff on travel)
  • May be ESO holiday

News / Meetings / Visitors

  • ALMA Dev. Study AstroPy: 7-18 April
    • No report.
  • ALMA ObsMode: April 14-16
    • Cycle III modes and capabilities under consideration.
    • Lots of desired features, more than likely possible, prioritizing may be necessary. Understood by Stewart and others.
    • Manual data reduction is not sustainable, need to get to 75% pipeline based data reduction.
    • Some give and take between what online system does vs data post processing system.
    • Trade-off between observing efficiency and data reduction effort is an active discussion. Also how much manual corrections to an SDM are acceptable.
    • Top priorities
      • Low SNR data reduction.
      • Observing efficiency.
      • Direct impact to CASA from priorities: Flagging will be impacted. Mostly pipeline affected. Some discussion of coordinate systems for Solar observing.
      • ACTION: Rob to ask Jeff to summarize ALMA Cycle 3 priorities and impact to CASA team.
    • ACTION: Lindsey follow up on ALMA correlator flags application.
  • Dirk Petry Visit: 12-23 May
  • Imaging Synthesis Workshop May 13-20
  • CASA Admin Calendar

Build, Release, Testing

  • CASA 4.2.1 Release on OSX
    • Punch list of automated tests to be resolved. Scott identifying which tests can be run headless. Will ask for assistance from team once they are identified.
    • Darrell & Susan on Viewer.
      • Test by Juergen confirmed not a packaging problem.
      • Scott setup Mac for Susan's development effort.
      • Next step is to merge trunk changes. Darrell is working on merging.
    • Mark will coordinate with other testers.
    • CASA 4.2.1. Plan
  • CASA 4.2.1 Release on EL
    • Public release concurrent with OSX release.
  • CASA 4.2.2
    • On hold last week in preference on 4.2.1 and 4.3 test package.
  • CASA 4.3 Test Packages
    • Scott delivered a test package end of Wednesday last week. Deployed on Thursday. (casapy test)


  • Technical Forums:
    • 5/1 is ESO holiday, no CTF.
    • Next will be Lindsey - Pipelines - May 15th
  • Modern compiler and C++ 11
    • No report
  • Environmental variables.
    • No report
  • CASA Document Repository.
    • No report


  • 4.2.2
    • Feature freeze on 4/25 (AKA Friday).
      • Most outstanding features will be ready.
      • Review progress next Monday and determine forward schedule.
    • CASA 4.2.2. Plan
    • Weights
      • George is close to committing weights changes on trunk. Then merge to the 4.2.2 branch.
      • Working on the regressions.SNR is impacted as expected. Need input from Dirk on key regression tests.
      • Application of bandpass solutions need some work.
      • Likely check in today or tomorrow, then on to merge list for 4.2.2.
    • George also trying to support Jason Koii while he is still here this week. Building animated display of TEC amongst other features.
    • ACTION: George will note changes needed from Michel for weights initialization in casacore on wiki page.
    • Flagging changes are included in the same commit George is working on. Will also be committed to trunk and eventually merged to 4.2.2
      • Existing behavior will not change, but you can enable this alternative stricter version of flagging.
    • Viewer
      • Concurrent to 4.2.1 development/debugging.
      • Non-reproducible crashes are an issue.
      • Need a dmg soon for testing. Mark to follow up.
    • Plotms iterations
      • Lindsey stresses it's important to include for processing time. Would strongly favor including this feature.
      • Susan was on testing and can likely support this feature. Will interweave this work along with 4.2.1 viewer work. Completion date TBD; review on Monday.
  • 4.3
    • Imaging - no report.
    • VO: progressing well. Integrating the viewer. Client is running and takes queries from a VO service running locally. Clean up work remains.
    • MPI - no report.


  • ALMA is focused on doing imaging with pipeline. Lots of pressure to get this feature implemented and generate images as a standard data product.
  • Self-calibration must be included to some level. Will require some consideration to implement reasonably.
  • Infrastructure development to support weblog meeting is still priority.
  • Implementation of plots are next priority (hence focus on plotms).
  • Improving reliability in the process.
  • Weblog user testing later this year.


  • iPython / sync - hold discussion next week.
  • Weekly staff reports - (halt practice?) - hold discussion next week.
  • Andy reported Mike's last day was last Friday. Coordinate any outstanding issues with Andy for the time being.
  • Mark to update JIRA ticket testing status today, expect e-mails.
Topic revision: r3 - 2014-04-21, RobSelina
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback